OpenStudy (anonymous):

So can someone help me solve this equation?? s = (-1/2*R*C) +/- ((1/2*R*C)^2 - (1/L*s))^1/2 . I am given values for R, C, and L. The problem being the s inside the parenthesis.

OpenStudy (anonymous):

@ybarrap , a stood for plus and minus in matlab.

OpenStudy (anonymous):

well it stood for [-1, 1] which performed the same task.

OpenStudy (anonymous):

Sorry but Im still so lost how -1/2rc becomes -RC/2

OpenStudy (ybarrap):

$$ s = \cfrac{-R C}{2} \pm \sqrt {\left (\cfrac{R C}{2}\right )^2 - \cfrac{s}{L}}\\ \implies\\ \cfrac{-R C}{2} + \sqrt {\left (\cfrac{R C}{2}\right )^2 - \cfrac{s}{L}}=\cfrac{-R C}{2} - \sqrt {\left (\cfrac{R C}{2}\right )^2 - \cfrac{s}{L}}\\ -\cfrac{R C}{2} + \sqrt {\left (\cfrac{R C}{2}\right )^2 - \cfrac{s}{L}}+ \cfrac{R C}{2} + \sqrt {\left (\cfrac{R C}{2}\right )^2 - \cfrac{s}{L}}=0\\ 2\sqrt {\left (\cfrac{R C}{2}\right )^2 - \cfrac{s}{L}}=0\\ \left (\cfrac{R C}{2}\right )^2 - \cfrac{s}{L}=0\\ \cfrac{s}{L}=\left (\cfrac{R C}{2}\right )^2\\ s=L\left (\cfrac{R C}{2}\right )^2\\ $$

OpenStudy (ybarrap):

Since you have that s= some number \(\pm\) something, we then know that some number \(+\) something=some number \(-\) something and so 2\(\times\)something=0.